OpenStudy (anonymous):

15. Explain how to solve the equation 3t + 11 = –10. Show that your solution works. (4 points)

OpenStudy (4n1m0s1ty):

3t + 11 (- 11) = -10 (- 11) 3t (/3) = -21 (/3) t = -7

OpenStudy (anonymous):

3t + 11 = -10 3t + 11 - 11 = -10 - 11 3t = -10 -11 3t = -21 3t/3 = -21/3 t = -7

OpenStudy (4n1m0s1ty):

Check 3*(-7) + 11 = -10 -21 + 11 = -10 -10 = -10

OpenStudy (anonymous):

Thank you so much i just didn't get this it's new to me i'm currently switching school and it's just been crazy i really Appreciate it... :D

OpenStudy (4n1m0s1ty):

np

OpenStudy (anonymous):

Try this one out. 5j + 17 = -13 Can you figure it out?

OpenStudy (anonymous):

j=4 ?

OpenStudy (anonymous):

Alright. What is the first step?

OpenStudy (anonymous):

See i'm sorry if i don't get it i have had math problems all of my life but umm what i did was 17 + -13 = 4 and i did not get it write can you please explain to me step by step?

OpenStudy (anonymous):

No prob. 1) Take 17 from both sides. 5j + 17 - 17 = -13 -17 What do we have now? Simplify.

OpenStudy (anonymous):

We now have: 5j = -13 - 17 We want to simplify this. 5j = -30 From here, we can divide both sides by whatever j is multiplied by to get 1 j. 5j/5 = j -30/5 = -6 So j = -6. Does this help?